OpenStudy (anonymous):

A tennis ball moves with a velocity of 25 m/s towards a wall. The ball bounces back with speed 22 m/s. What was the average acceleration of the ball if the bounce lasted 3.2 ms?

OpenStudy (ksaimouli):

3.2 time or velocity?

OpenStudy (anonymous):

velocity, milliseconds

OpenStudy (ksaimouli):

convert that ms in to m/s

OpenStudy (ksaimouli):

initial velocity =25 and final vel;ocity is 22

OpenStudy (anonymous):

velocity = 22m/s time = 3,2ms = 0,0032 seconds formula: a = v/t so: 22/0,0032 = 6875 except the answer is 15000m/s^2

OpenStudy (anonymous):

hmm... I think I'm misundertanding the question

OpenStudy (anonymous):

delta V = Vf -Vi = 22 - -25 = 47 m/s 47 m/s /(.0032s) = 14,687m/s^2

OpenStudy (anonymous):

\[a=\frac{ \Delta V }{ t }\]